A monatomic ideal gas at 27°C undergoes an isobaric process from state A to B, followed by an isochoric process from state B to C. What is the total work done by the gas in these two processes? (A = 2 atm, 1 L; B = 2 atm, 2 L; C = 1 atm, 2 L)
1
Expert's answer
2020-10-25T18:50:12-0400

"T = 27\degree{C} = 300K"


"A = 2atm, \ 1L\ (i.e \ p_1 =2\,atm,\ V_1 =1L )\\
B = 2atm, \ 2L\ (i.e \ p_2 =2\,atm,\ V_2 =2L )\\
C = 1atm, \ 2L\ (i.e \ p_3 =1\,atm,\ V_1 =2L )"


"Process_{A\to B}=Isobaric\\
Process_{B\to C}=Isochoric"


"Work\ during\ an\ isobaric\ process=p(V_2-V_1)\\
W_{A\to B} = p(V_2-V_1)\\
W_{A\to B} = 2(2-1)\\
W_{A\to B} = 2(1)\\
W_{A\to B} = 2\,atmL"


"Work\ during\ an\ isochoric\ process=0J\\
W_{B\to C} = 0J = 0\,atmL"

(This is because the volume is constant, hence no work is done on the surroundings)


"\begin{aligned}

Total\ Work\ Done& = W_{A\to B} + W_{B\to C}\\
&= 2\,atmL + 0\,atmL\\
& = 2\,atmL\\
&= 2(101325Nm^2/atm)(10^{-3}m^3/L)atmL\\
&= 202.65J

\end{aligned}"


"\therefore" The Total Work-done by the gas in the two processes is "202.65J"